Make the “line of the best fit” for the number of hours you used your cell phone in a week. Let your day represent the y - axis and the number of hours you us your cellphone represent the x – axis.


1
Expert's answer
2021-10-24T18:29:39-0400

Day 1 - Number of Hours: 3

Day 2 - Number of Hours: 4

Day 3 - Number of Hours: 2.5

Day 4 - Number of Hours: 4.2

Day 5 - Number of Hours: 3.5

Day 6 - Number of Hours: 2.8

Day 7 - Number of Hours: 4


line of the best fit:

y=ax+b


"a=\frac{\sum y\sum x^2-\sum x\sum xy}{n\sum x^2-(\sum x)^2}"


"b=\frac{n\sum xy-\sum x\sum y}{n\sum x^2-(\sum x)^2}"


where x is number of day,

y is the number of hours


"y=0.057x+3.2"